OpenStudy (anonymous):

Simplify the following expression. a + 2c - 5b - c + a - b

OpenStudy (callisto):

a + 2c - 5b - c + a - b = a + a - 5b - b + 2c - c = ...?

OpenStudy (anonymous):

???

OpenStudy (callisto):

Can you do it from there? Combine the like terms..

OpenStudy (anonymous):

im really stuck on it

OpenStudy (callisto):

Example: a+a = (1+1) a = 2a What is -5b-b?

OpenStudy (anonymous):

-6b ?

OpenStudy (callisto):

Yes. Now, what is +2c-c?

OpenStudy (anonymous):

2a - 6b + c i got it (:

OpenStudy (callisto):

Cool!!~ See? You're NOT stuck on it :)

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

Welcome :)